# Tensor global symmetries and the Stueckelberg mechanism for tensor fields

### v1

@@ 4.2 Tensor global symmetries & gauging
$S[a,\phi]=\int\mathrm{d}^{4}x\big{(}-\frac{1}{2}\partial_{\mu}a_{\nu}\partial^{\mu}a^{\nu}+\frac{1}{2}\partial^{\mu}a_{\mu}\partial^{\nu}a_{\nu}+\,\Box\phi\,\Box\phi-\partial_{\mu}\partial_{\nu}\phi\,\partial^{\mu}\partial^{\nu}\phi\big{)}\,.$ (40)
@@
$J^{(a)}_{\mu\nu}=\frac{1}{2}(\partial_{\mu}a_{\nu}+\partial_{\nu}a_{\mu})-\eta_{\mu\nu}\partial^{\rho}a_{\rho}+\frac{1}{2}F_{\mu\nu}\,.$ (41)
@@
$J^{(a)}_{(\mu\nu)}$ $=\frac{1}{2}(\partial_{\mu}a_{\nu}+\partial_{\nu}a_{\mu})-\eta_{\mu\nu}\partial^{\rho}a_{\rho},$ (43)
$J^{(a)}_{[\mu\nu]}$ $=\frac{1}{2}F_{\mu\nu}\,,$ (44)
@@
$J^{\text{e}}_{\mu\nu}H^{\mu\nu}=\partial_{\mu}a_{\nu}H^{\mu\nu}-\partial^{\rho}a_{\rho}H^{\mu}{}_{\mu}+\,\partial_{\mu}\partial_{\nu}\phi H^{\mu\nu}-\Box\phi H^{\mu}{}_{\mu}\,.$ (48)
These couplings are precisely the ones we found in the Stueckelberg approach, Eq. (37), which justifies that this is the right current.
@@
$J^{(a)}_{\mu}{}^{\mu}=-3\,\partial^{\mu}a_{\mu}\quad\text{and}\quad J^{(\phi)}_{\mu}{}^{\mu}=-3\,\Box\phi\,,$ (49)
which appear in the minimal coupling. This also shows that the status of $a$ and $\phi$ is different, as already seen through Weyl rescaling (38). The trace of $J^{(a)}$ is the divergence of $a_{\mu}$ that appears in the Lorenz gauge condition, whereas the trace of $J^{(\phi)}$ is proportional to the d’Alembertian of $\phi$ that appears in the wave equation.
Gauging is performed in the usual way. Adding the minimal coupling (48) to the original Lagrangian (40), a “seagull” term quadratic in $H_{\mu\nu}$ makes the kinetic term invariant under the shift symmetry with parameter $\alpha_{\mu}$. Then we may add the kinetic term for $H_{\mu\nu}$, turning the background field into a dynamical one. It is worth recalling that in the massless limit, the theory at this level does not result in the linearized limit of general relativity (vDVZ discontinuity) [41, 42], unless higher-order terms in $H_{\mu\nu}$ are included [43].
@@ 4.3 Alternative view of the Noether symmetry
There is an equivalent yet seemingly different view of the symmetry transformations (29)-(31). We can rewrite them as
@@
$\delta a_{\mu}$ $=\alpha_{\mu}-\partial_{\mu}{\varphi}\,,$ (50)
@@
This form is more natural from the Noether perspective. Firstly, we do not encounter a two-derivative transformation (29) and secondly the gauge transformation $\delta a_{\mu}=-\partial_{\mu}{\varphi}$ is manifest and hence the gauge invariance of the current seems more natural. Moreover, looking at the symmetries in this manner, the auxiliary symmetric current (45) is nothing but the freedom in the definition of the Noether current which is unique up to an exact term [44]. It means one can improve the symmetric conserved Noether current $J^{(a)}_{(\mu\nu)}$ by a divergence term, i.e.,
$J^{\text{e}}_{\mu\nu}=J^{(a)}_{(\mu\nu)}+\partial^{\rho}k^{(\phi)}_{\rho\mu\nu}\,,$ (51)
where $k^{\rho\mu\nu}=k^{\rho(\mu\nu)}$ and $k^{(\rho\mu\nu)}=0$, subject to the constraint that $\partial^{\mu}J^{\text{e}}_{\mu\nu}=0$. From Eq. (45), we can see that
$k^{\rho\mu\nu}=\eta^{\rho(\mu}\partial^{\nu)}\phi-\eta^{\mu\nu}\partial^{\rho}\phi\,,$ (52)
@@
There is an important distinction between the Stueckelberg field $a_{\mu}$ and the “secondary” Stueckelberg field $\phi$, that can be understood from their transformation under the shift symmetry (4.3). The “secondary” Stueckelberg field $\phi$ cannot become the Goldstone boson. This is due to the fact that its global shift symmetry has already been gauged by $a_{\mu}$. In fact, this is a general behavior; only those Stueckelberg fields can become Goldstone bosons as a result of symmetry breaking if in the corresponding set of symmetry transformations as in (4.3) their shift symmetry is gauged by the background field and not by another Stueckelberg field. That is to say, the terminology “secondary” Stueckelberg field indicates that it cannot become a Goldstone boson. Another way to understand this point is by noticing that after performing the Weyl transformation (38), there will appear a mass term for the “secondary” Stueckelberg field contrary to the other Stueckelberg field that is massless. It is well understood that only massless fields can become Goldstone bosons [45].
@@ 5.1 Stueckelberg action for the massive Curtright field
$T_{(\mu\nu)|\rho}=0\quad\text{and}\quad T_{[\mu\nu|\rho]}=0\,.$ (53)
@@
$\hskip 28.0pt=-\frac{1}{2}\int\mathrm{d}^{4}x\int_{\text{B}}\bigg{(}\mathrm{d}T\star\mathrm{d}T-T\star T\bigg{)}\,.$ (54)
@@
$\delta T_{\mu\nu|\rho}=2\partial_{[\mu}s_{\nu]\rho}+2\partial_{[\mu}\beta_{\nu]\rho}-2\partial_{\rho}\beta_{\mu\nu}\,,$ (55)
@@
$\delta T_{\mu\nu|\rho}=2\partial_{\rho}\partial_{[\mu}\alpha_{\nu]}\,,$ (56)
@@
$\delta h_{\mu\nu}=s_{\mu\nu}\,,\quad\delta b_{\mu\nu}=\beta_{\mu\nu}\,.\quad\delta a_{\mu}=\alpha_{\mu}\,.$ (57)
@@
$T_{\mu\nu|\rho}\to\mathring{F}_{\mu\nu|\rho}:=T_{\mu\nu|\rho}-2\partial_{[\mu}h_{\nu]\rho}-2\partial_{[\mu}b_{\nu]\rho}+2\partial_{\rho}b_{\mu\nu}-2\partial_{\rho}\partial_{[\mu}a_{\nu]}\,.$ (58)
@@
$\mathring{F}=T-\mathrm{d}h-\mathrm{d}\widetilde{\sigma}b+2\widetilde{\mathrm{d}}b-2\mathrm{d}\widetilde{\mathrm{d}}a\,,$ (60)
@@
$\widetilde{\sigma}=-\widetilde{\theta}^{\mu}\frac{\partial}{\partial\theta^{\mu}}\,,$ (61)
@@
$S_{\text{Curt}}^{\text{St}}=$ $-\frac{1}{2}\int\mathrm{d}^{4}x\int_{\text{B}}(\mathrm{d}T-2\mathrm{d}\widetilde{\mathrm{d}}b)\star(\mathrm{d}T-2\mathrm{d}\widetilde{\mathrm{d}}b)$ (62)
@@
$S_{\text{Curt}}^{\text{St}}$ $=$ $-\frac{1}{2}\int\mathrm{d}^{4}x\int_{\text{B}}\big{(}\mathrm{d}T\star\mathrm{d}T-T\star T\big{)}$ (63)
@@
$-T_{\mu\nu|}{}^{\mu}(\partial^{\nu}\partial^{\rho}a_{\rho}-\Box a^{\nu})-T_{\mu\nu|\rho}\partial^{\rho}\partial^{\mu}a^{\nu}\bigg{)}\,,$ (64)
@@
$H_{\mu\nu\rho}=\partial_{\mu}b_{\nu\rho}+\partial_{\rho}b_{\mu\nu}+\partial_{\nu}b_{\rho\mu}\,,$ (65)
@@
$\qquad+\frac{1}{8}\int\mathrm{d}^{4}x\bigg{(}(\Box a_{\nu}-\partial_{\nu}\partial_{\mu}a^{\mu})(\Box a^{\nu}-\partial^{\nu}\partial^{\rho}a_{\rho})-\,\partial_{\mu}\partial_{\nu}a_{\rho}\partial^{\nu}(\partial^{\rho}a^{\mu}-\partial^{\mu}a^{\rho})\bigg{)}\,.$ (67)
The first term is just the usual linearized Einstein-Hilbert action, given by (23) without the mass terms. The second term is an unorthodox way to rewrite the action for a Kalb-Ramond field, essentially after performing an integration by parts and dropping the boundary term. The sector that contains the 1-form $a_{\mu}$ is also unorthodox, it is just a complicated way to write a total derivative, and it does not contribute to the field equations.
@@
$J_{\mu\nu|\rho}^{(h)}$ $=$ $\partial_{[\mu}h_{\nu]\rho}+\eta_{\rho[\mu}\partial_{\nu]}h_{\sigma}{}^{\sigma}-\eta_{\rho[\mu}\partial^{\sigma}h_{\nu]\sigma}\,,$ (68)
$J_{\mu\nu\rho}^{(b)}$ $=$ $\partial_{\rho}b_{\mu\nu}+\partial_{[\mu}b_{\nu]\rho}+\eta_{\rho[\mu}\partial^{\sigma}b_{\nu]\sigma}\,,$ (69)
$J_{\mu\nu|\rho}^{(a)}$ $=$ $\partial_{\rho}\partial_{[\mu}a_{\nu]}+\eta_{\rho[\mu}\partial_{\nu]}\partial^{\sigma}a_{\sigma}-\eta_{\rho[\mu}\Box a_{\nu]}\,.$ (70)
$J^{(h)}$ and $J^{(a)}$ are genuine $(2,1)$ tensors, since they are antisymmetric in the first two indices by construction and their fully antisymmetric part vanishes. $J^{(b)}$ on the other hand, has a fully antisymmetric part (note the absence of vertical index separator in (69)), which would anyway vanish when contracted with an irreducible $(2,1)$ tensor. In fact, a genuine $(2,1)$ current for the Kalb-Ramond field may be defined by noting that $J^{(b)}_{[\mu\nu\rho]}=\frac{2}{3}H_{\mu\nu\rho}$. Then we can subtract the fully antisymmetric part from (69) and construct a genuine $(2,1)$ current that reads:
$J_{\mu\nu|\rho}^{(b)}=\frac{1}{3}\big{(}\partial_{\rho}b_{\mu\nu}-\partial_{[\mu}b_{\nu]\rho}+3\eta_{\rho[\mu}\partial^{\sigma}b_{\nu]\sigma}\big{)}\,.$ (71)
@@
$\partial^{\mu}J^{(h)}_{\mu\nu|\rho}$ $=$ $\frac{1}{2}(\Box h_{\nu\rho}-\partial^{\mu}\partial_{\nu}h_{\mu\rho}+\eta_{\rho\mu}\partial^{\mu}\partial_{\nu}h_{\sigma}{}^{\sigma}$ (73)
@@
$\partial^{\rho}J^{(b)}_{\mu\nu|\rho}=\partial^{\rho}J^{(b)}_{\rho\mu|\nu}=\frac{1}{3}\partial^{\rho}H_{\mu\nu\rho}\,,$ (74)
@@
$\delta a_{\mu}$ $=$ $-\alpha_{\mu}+\partial_{\mu}\varepsilon\,,$ (75)
$\delta b_{\mu\nu}$ $=$ $2\partial_{[\mu}\alpha_{\nu]}\,,$ (76)
$\delta h_{\mu\nu}$ $=$ $2\partial_{(\mu}\alpha_{\nu)}+\partial_{\mu}\partial_{\nu}\varepsilon\,.$ (77)
@@
$\delta J^{(h)}=\delta J^{(b)}=-\delta J^{(a)}=\partial_{\rho}\partial_{[\mu}\alpha_{\nu]}+\eta_{\rho[\mu}\partial_{\nu]}\partial^{\sigma}\alpha_{\sigma}-\eta_{\rho[\mu}\Box\alpha_{\nu]}\,.$ (78)
@@
$J^{(h)}_{\mu\nu}{}^{\mu}$ $=$ $\partial_{\nu}h_{\mu}{}^{\mu}-\partial^{\mu}h_{\mu\nu}\,,$ (81)
$J^{(b)}_{\mu\nu}{}^{\mu}$ $=$  $-\,2\partial^{\mu}b_{\mu\nu}\,,$ (82)
$J^{(a)}_{\mu\nu}{}^{\mu}$ $=$ $-\,\partial^{\mu}F_{\mu\nu}\,.$ (83)
@@
It is instructive to compare our results to previous approaches on the graviton as a Nambu-Goldstone boson. The conserved and gauge-invariant current (79), when the Kalb-Ramond and the vector fields are set to zero, reduces to the current (68), which was calculated in [26] as
$J^{({h})}_{\mu\nu|}{}^{\rho}=3\delta_{\mu\nu\sigma}^{\kappa\lambda\rho}\partial_{[\kappa}h_{\lambda]}^{\sigma}\,,$ (84)
where $\delta_{\mu\nu\sigma}^{\kappa\lambda\rho}$ is the generalized Kronecker delta. However, from the standpoint of our approach, this does not constitute a consistent truncation, as the matching of physical polarizations in the Stueckelberg mechanism would fail. Furthermore, the current (84) is not gauge-invariant under the shift symmetry of $h_{\mu\nu}$, i.e., (77) with $\varepsilon=0$.
@@
$\delta h_{\mu\nu}=2\partial^{\rho}\Lambda_{\rho(\mu|\nu)}\,,$ (85)

### v2

@@ 4.2 Tensor global symmetries & gauging
$S[a,\phi]=\int\mathrm{d}^{4}x\big{(}-\frac{1}{2}\partial_{\mu}a_{\nu}\partial^{\mu}a^{\nu}+\frac{1}{2}\partial^{\mu}a_{\mu}\partial^{\nu}a_{\nu}+\,\Box\phi\,\Box\phi-\partial_{\mu}\partial_{\nu}\phi\,\partial^{\mu}\partial^{\nu}\phi\big{)}\,.$ (4.18)
@@
$J^{(a)}_{\mu\nu}=\frac{1}{2}(\partial_{\mu}a_{\nu}+\partial_{\nu}a_{\mu})-\eta_{\mu\nu}\partial^{\rho}a_{\rho}+\frac{1}{2}F_{\mu\nu}\,.$ (4.19)
@@
$J^{(a)}_{(\mu\nu)}$ $=\frac{1}{2}(\partial_{\mu}a_{\nu}+\partial_{\nu}a_{\mu})-\eta_{\mu\nu}\partial^{\rho}a_{\rho},$ (4.21)
$J^{(a)}_{[\mu\nu]}$ $=\frac{1}{2}F_{\mu\nu}\,,$ (4.22)
@@
$J^{\text{e}}_{\mu\nu}H^{\mu\nu}=\partial_{\mu}a_{\nu}H^{\mu\nu}-\partial^{\rho}a_{\rho}H^{\mu}{}_{\mu}+\,\partial_{\mu}\partial_{\nu}\phi H^{\mu\nu}-\Box\phi H^{\mu}{}_{\mu}\,.$ (4.26)
These couplings are precisely the ones we found in the Stueckelberg approach, Eq. (4.15), which justifies that this is the right current.
@@
$J^{(a)}_{\mu}{}^{\mu}=-3\,\partial^{\mu}a_{\mu}\quad\text{and}\quad J^{(\phi)}_{\mu}{}^{\mu}=-3\,\Box\phi\,,$ (4.27)
which appear in the minimal coupling. This also shows that the status of $a$ and $\phi$ is different, as already seen through Weyl rescaling (4.16). The trace of $J^{(a)}$ is the divergence of $a_{\mu}$ that appears in the Lorenz gauge condition, whereas the trace of $J^{(\phi)}$ is proportional to the d’Alembertian of $\phi$ that appears in the wave equation.
Gauging is performed in the usual way. Adding the minimal coupling (4.26) to the original Lagrangian (4.18), a “seagull” term quadratic in $H_{\mu\nu}$ makes the kinetic term invariant under the shift symmetry with parameter $\alpha_{\mu}$. Then we may add the kinetic term for $H_{\mu\nu}$, turning the background field into a dynamical one. It is worth recalling that in the massless limit, the theory at this level does not result in the linearized limit of general relativity (vDVZ discontinuity) [50, 51], unless higher-order terms in $H_{\mu\nu}$ are included [52].
@@ 4.3 Alternative view of the Noether symmetry
There is an equivalent yet seemingly different view of the symmetry transformations (4.7)-(4.9). We can rewrite them as
@@
$\delta a_{\mu}$ $=\alpha_{\mu}-\partial_{\mu}{\varphi}\,,$ (4.28)
@@
This form is more natural from the Noether perspective. Firstly, we do not encounter a two-derivative transformation (4.7) and secondly the gauge transformation $\delta a_{\mu}=-\partial_{\mu}{\varphi}$ is manifest and hence the gauge invariance of the current seems more natural. Moreover, looking at the symmetries in this manner, the auxiliary symmetric current (4.23) is nothing but the freedom in the definition of the Noether current which is unique up to an exact term [53]. It means one can improve the symmetric conserved Noether current $J^{(a)}_{(\mu\nu)}$ by a divergence term, i.e.,
$J^{\text{e}}_{\mu\nu}=J^{(a)}_{(\mu\nu)}+\partial^{\rho}k^{(\phi)}_{\rho\mu\nu}\,,$ (4.29)
where $k^{\rho\mu\nu}=k^{\rho(\mu\nu)}$ and $k^{(\rho\mu\nu)}=0$, subject to the constraint that $\partial^{\mu}J^{\text{e}}_{\mu\nu}=0$. From Eq. (4.23), we can see that
$k^{\rho\mu\nu}=\eta^{\rho(\mu}\partial^{\nu)}\phi-\eta^{\mu\nu}\partial^{\rho}\phi\,,$ (4.30)
@@
There is an important distinction between the Stueckelberg field $a_{\mu}$ and the “secondary” Stueckelberg field $\phi$, that can be understood from their transformation under the shift symmetry (4.3). The “secondary” Stueckelberg field $\phi$ cannot become the Goldstone boson. This is due to the fact that its global shift symmetry has already been gauged by $a_{\mu}$. In fact, this is a general behavior; only those Stueckelberg fields can become Goldstone bosons as a result of symmetry breaking if in the corresponding set of symmetry transformations as in (4.3) their shift symmetry is gauged by the background field and not by another Stueckelberg field. That is to say, the terminology “secondary” Stueckelberg field indicates that it cannot become a Goldstone boson. Another way to understand this point is by noticing that after performing the Weyl transformation (4.16), there will appear a mass term for the “secondary” Stueckelberg field contrary to the other Stueckelberg field that is massless. It is well understood that only massless fields can become Goldstone bosons [54].
@@ 5.1 Stueckelberg action for the massive Curtright field
$T_{(\mu\nu)|\rho}=0\quad\text{and}\quad T_{[\mu\nu|\rho]}=0\,.$ (5.1)
@@
$\hskip 28.0pt=-\frac{1}{2}\int\mathrm{d}^{4}x\int_{\text{B}}\bigg{(}\mathrm{d}T\star\mathrm{d}T-T\star T\bigg{)}\,.$ (5.2)
@@
$\delta T_{\mu\nu|\rho}=2\partial_{[\mu}s_{\nu]\rho}+2\partial_{[\mu}\beta_{\nu]\rho}-2\partial_{\rho}\beta_{\mu\nu}\,,$ (5.3)
@@
$\delta T_{\mu\nu|\rho}=2\partial_{\rho}\partial_{[\mu}\alpha_{\nu]}\,,$ (5.4)
@@
$\delta h_{\mu\nu}=s_{\mu\nu}\,,\quad\delta b_{\mu\nu}=\beta_{\mu\nu}\,.\quad\delta a_{\mu}=\alpha_{\mu}\,.$ (5.5)
@@
$T_{\mu\nu|\rho}\to\mathring{F}_{\mu\nu|\rho}:=T_{\mu\nu|\rho}-2\partial_{[\mu}h_{\nu]\rho}-2\partial_{[\mu}b_{\nu]\rho}+2\partial_{\rho}b_{\mu\nu}-2\partial_{\rho}\partial_{[\mu}a_{\nu]}\,.$ (5.6)
@@
$\mathring{F}=T-\mathrm{d}h-\mathrm{d}\widetilde{\sigma}b+2\widetilde{\mathrm{d}}b-2\mathrm{d}\widetilde{\mathrm{d}}a\,,$ (5.8)
@@
$\widetilde{\sigma}=-\widetilde{\theta}^{\mu}\frac{\partial}{\partial\theta^{\mu}}\,,$ (5.9)
@@
$S_{\text{Curt}}^{\text{St}}=$ $-\frac{1}{2}\int\mathrm{d}^{4}x\int_{\text{B}}(\mathrm{d}T-2\mathrm{d}\widetilde{\mathrm{d}}b)\star(\mathrm{d}T-2\mathrm{d}\widetilde{\mathrm{d}}b)$ (5.10)
@@
$S_{\text{Curt}}^{\text{St}}$ $=$ $-\frac{1}{2}\int\mathrm{d}^{4}x\int_{\text{B}}\big{(}\mathrm{d}T\star\mathrm{d}T-T\star T\big{)}$ (5.11)
@@
$-T_{\mu\nu|}{}^{\mu}(\partial^{\nu}\partial^{\rho}a_{\rho}-\Box a^{\nu})-T_{\mu\nu|\rho}\partial^{\rho}\partial^{\mu}a^{\nu}\bigg{)}\,,$ (5.12)
@@
$H_{\mu\nu\rho}=\partial_{\mu}b_{\nu\rho}+\partial_{\rho}b_{\mu\nu}+\partial_{\nu}b_{\rho\mu}\,,$ (5.13)
@@
$\qquad+\frac{1}{8}\int\mathrm{d}^{4}x\bigg{(}(\Box a_{\nu}-\partial_{\nu}\partial_{\mu}a^{\mu})(\Box a^{\nu}-\partial^{\nu}\partial^{\rho}a_{\rho})-\,\partial_{\mu}\partial_{\nu}a_{\rho}\partial^{\nu}(\partial^{\rho}a^{\mu}-\partial^{\mu}a^{\rho})\bigg{)}\,.$ (5.15)
The first term is just the usual linearized Einstein-Hilbert action, given by (4.1) without the mass terms. The second term is an unorthodox way to rewrite the action for a Kalb-Ramond field, essentially after performing an integration by parts and dropping the boundary term. The sector that contains the 1-form $a_{\mu}$ is also unorthodox, it is just a complicated way to write a total derivative, and it does not contribute to the field equations.
@@
$J_{\mu\nu|\rho}^{(h)}$ $=$ $\partial_{[\mu}h_{\nu]\rho}+\eta_{\rho[\mu}\partial_{\nu]}h_{\sigma}{}^{\sigma}-\eta_{\rho[\mu}\partial^{\sigma}h_{\nu]\sigma}\,,$ (5.16)
$J_{\mu\nu\rho}^{(b)}$ $=$ $\partial_{\rho}b_{\mu\nu}+\partial_{[\mu}b_{\nu]\rho}+\eta_{\rho[\mu}\partial^{\sigma}b_{\nu]\sigma}\,,$ (5.17)
$J_{\mu\nu|\rho}^{(a)}$ $=$ $\partial_{\rho}\partial_{[\mu}a_{\nu]}+\eta_{\rho[\mu}\partial_{\nu]}\partial^{\sigma}a_{\sigma}-\eta_{\rho[\mu}\Box a_{\nu]}\,.$ (5.18)
$J^{(h)}$ and $J^{(a)}$ are genuine $(2,1)$ tensors, since they are antisymmetric in the first two indices by construction and their fully antisymmetric part vanishes. $J^{(b)}$ on the other hand, has a fully antisymmetric part (note the absence of vertical index separator in (5.17)), which would anyway vanish when contracted with an irreducible $(2,1)$ tensor. In fact, a genuine $(2,1)$ current for the Kalb-Ramond field may be defined by noting that $J^{(b)}_{[\mu\nu\rho]}=\frac{2}{3}H_{\mu\nu\rho}$. Then we can subtract the fully antisymmetric part from (5.17) and construct a genuine $(2,1)$ current that reads:
$J_{\mu\nu|\rho}^{(b)}=\frac{1}{3}\big{(}\partial_{\rho}b_{\mu\nu}-\partial_{[\mu}b_{\nu]\rho}+3\eta_{\rho[\mu}\partial^{\sigma}b_{\nu]\sigma}\big{)}\,.$ (5.19)
@@
$\partial^{\mu}J^{(h)}_{\mu\nu|\rho}$ $=$ $\frac{1}{2}(\Box h_{\nu\rho}-\partial^{\mu}\partial_{\nu}h_{\mu\rho}+\eta_{\rho\mu}\partial^{\mu}\partial_{\nu}h_{\sigma}{}^{\sigma}$ (5.21)
@@
$\partial^{\rho}J^{(b)}_{\mu\nu|\rho}=\partial^{\rho}J^{(b)}_{\rho\mu|\nu}=\frac{1}{3}\partial^{\rho}H_{\mu\nu\rho}\,,$ (5.22)
@@
$\delta a_{\mu}$ $=$ $-\alpha_{\mu}+\partial_{\mu}\varepsilon\,,$ (5.23)
$\delta b_{\mu\nu}$ $=$ $2\partial_{[\mu}\alpha_{\nu]}\,,$ (5.24)
$\delta h_{\mu\nu}$ $=$ $2\partial_{(\mu}\alpha_{\nu)}+\partial_{\mu}\partial_{\nu}\varepsilon\,.$ (5.25)
@@
$\delta J^{(h)}=\delta J^{(b)}=-\delta J^{(a)}=\partial_{\rho}\partial_{[\mu}\alpha_{\nu]}+\eta_{\rho[\mu}\partial_{\nu]}\partial^{\sigma}\alpha_{\sigma}-\eta_{\rho[\mu}\Box\alpha_{\nu]}\,.$ (5.26)
@@
$J^{(h)}_{\mu\nu}{}^{\mu}$ $=$ $\partial_{\nu}h_{\mu}{}^{\mu}-\partial^{\mu}h_{\mu\nu}\,,$ (5.29)
$J^{(b)}_{\mu\nu}{}^{\mu}$ $=$  $-\,2\partial^{\mu}b_{\mu\nu}\,,$ (5.30)
$J^{(a)}_{\mu\nu}{}^{\mu}$ $=$ $-\,\partial^{\mu}F_{\mu\nu}\,.$ (5.31)
@@
It is instructive to compare our results to previous approaches on the graviton as a Nambu-Goldstone boson. The conserved and gauge-invariant current (5.27), when the Kalb-Ramond and the vector fields are set to zero, reduces to the current (5.16), which was calculated in [28] as
$J^{({h})}_{\mu\nu|}{}^{\rho}=3\delta_{\mu\nu\sigma}^{\kappa\lambda\rho}\partial_{[\kappa}h_{\lambda]}^{\sigma}\,,$ (5.32)
where $\delta_{\mu\nu\sigma}^{\kappa\lambda\rho}$ is the generalized Kronecker delta. However, from the standpoint of our approach, this does not constitute a consistent truncation, as the matching of physical polarizations in the Stueckelberg mechanism would fail. Furthermore, the current (5.32) is not gauge-invariant under the shift symmetry of $h_{\mu\nu}$, i.e., (5.25) with $\varepsilon=0$.
@@
$\delta h_{\mu\nu}=2\partial^{\rho}\Lambda_{\rho(\mu|\nu)}\,,$ (5.33)